# Hyperparameter Optimization in Machine Learning

### v2

@@ 8.5 Transfer of hyperparameters across model scale
In this chapter, we provide an overview of different open-source libraries and commercial systems that implement methods discussed in the previous chapters. HPO software landscape changes frequently as new problems, methods and application domain become relevant: the provided list here is by no means exhaustive and only represents a snapshot.
@@ 9.1 Open Source Libraries and Frameworks
While previous libraries, such as Spearmint, allowed to parallelize trials across processes on the same machine, more recent libraries focus more on the large-scale distributed setting to distribute the evaluation of trials across multiple instances or compute nodes. For example, RayTune [liaw2018tune], distributes the evaluation of hyperparameter configurations using the compute framework Ray. Another example is Orion, which supports both high-performance computing infrastructures and containerized environments. salinas-automl22 recently presented Syne Tune, which supports different backends that allow to distribute the HPO process either locally or on the cloud. Furthermore, it also provides a backend to simulate the evaluation of expensive objective functions using surrogate models (see Section 9.3). This allows to simulate large-scale asynchronous HPO on a single compute core.
@@
Most HPO libraries or frameworks expect the objective function either as function handle, for example in Python, or as a single training script. The search space is usually defined outside of the actual training script and hyperparameters are passed as input arguments. Optuna [akiba-sigkdd19] follows a different approach by a so-called define-by-run API that allows users to dynamically define the search space instead of providing it in advance.
@@ 9.2 Commercial Systems
Since hyperparameter optimization plays such an integral role in the machine learning workflow, most machine learning platforms, for example Weights and Biases, provide some functionality to select hyperparameters. There are also several commercial systems specific for automated hyperparameter optimization. For example, Google Vizier [google_vizier] is an internal service that contributed to several Google products and research efforts. oss_vizier also introduces a open-source version of Google Vizier as a stand-alone Python package. AWS SageMaker provides a fully automated service called Automated Model Tuner [perrone2021sagemaker] for gradient-free optimization problems, which allows for large-scale HPO leveraging built-in SageMaker functionalities. Another example is Sigopt, which provides optimization-as-a-service for hyperparameter optimization utilizing Bayesian optimization techniques.
@@ 9.3 Benchmarking
Instead of using a machine learning model to approximate the true objective, klein-arxiv19a first discretzied the search space of a fully-connected neural network leading to a finite number of hyperparameter configurations and then performed an exhaustive search to collect the performance and runtime of each configuration. This allows to replace the actual training and validation of a hyperparameter configuration by simple table look-ups during the HPO process. To mimic the observation noise that comes with the HPO of neural networks, each HPO configuration was evaluated multiple times. During the HPO process, random entries for each hyperparameter configuration are returned to simulate the observation noise. The idea of tabularized benchmarks has since been extended to neural architecture search [ying-icml19, dong-iclr20] (see Section 8.3) leading to several new benchmarks. While tabularized benchmarks return for each query truly observed values of the objective function instead of prediction of the model, pfisterer-automl22 argued that the discretization of the search space eliminates the local structure of the objective function, leading actually to a higher bias than using surrogate models. The idea of surrogate benchmarks has been extended to define a generative model by [klein-nips19], which allows sampling an arbitrary set of new benchmarks.
To standardize the interface to different benchmarks, eggensperger_towards_2013 introduced HPOLib, which consolidates different benchmarks from the literature. In follow up work [eggensperger-neurips21], HPOBench containerized these benchmarks to increase reproducibility of benchmarking results. HPO-B [pineda-neurips21] provides tablularized benchmarks based on OpenML datasets [vanschoren-sigkdd13a]. In a similar vein, zimmer-ieee21 introduced LCBench, which provides full learning curves of neural network models trained on OpenML datasets. salinas-automl22 also provided a unified interface to tabular and surrogate benchmarks in Syne-Tune, using a fast data format to accelerate the table look-ups. Another suite for surrogate-based benchmarking is YAHPO [pfisterer-automl22], consisting of different scenarios, including combined hyperparameter optimization and algorithm selection.
@@
From some perspective, hyperparameter optimization may be considered by now a mature field with well-established algorithmic solutions and with industry-level software frameworks. It is a crucial ingredient for automated machine learning, an active research area that has even started a conference in 2022. In this monograph, we have reviewed in depth the major classes of methods (random and quasi-random search, model-based, multi-fidelity, population-based, and gradient-based) and highlighted connections to closely related areas such neural architecture search and meta-learning.
@@ 10.1 Response functions for unsupervised learning
HPO requires a well defined response function to be minimized (see Section 2.2), which in turn requires a well defined performance metric. While suitable metrics are abundant for supervised learning, the same is not true for unsupervised learning.
@@
There have been interesting attempts to interpret this perspective in terms of precision and recall [sajjadi:2018:assessing_generative_models, simon:2019:revisiting_precision_recall], where the two supports (of real and fake data) are estimated and compared (fake points off the support of real distribution are interpreted as false positives, and real points outside the support of the model’s distribution are interpreted as false negatives). The idea has sparkled a number of practically computable metrics [kynkaanniemi:2019:improved_precision_recall, naeem:2020:reliable_fidelity_diversitya, park:2023:probabilistic_precision_recall] that however have not been shown to be effective to guide HPO. A unidimensional metric that aims to assess the similarity between the model and the real data distributions is the overlapping area. Computing the overlaps between these two distributions requires reliable density estimation, which is however difficult for high dimensional data. A simple trick to overcome the associated curse of dimensionalty issue is to consider instead the distribution of intra-set distances (between pairs of real data points) and inter-set distances (between pairs of fake points). This metric was introduced in [yang:2020:evaluation_generative_models] for assessing the quality of generative models in symbolic music, but distances in that case are based on domain specific aspects of music such as rhythm or harmony, and cannot be easily extended to other types of data. A closely related metric based on maximum mean discrepancy [obray:2021:evaluation_metrics_graph], popular in the context of graph generative models, has been shown to be complicated due to its strong sensitivity to hyperparameters in the definition of the metric itself (leading to a sort of hyper-hyperparameters).
@@
10.2 Learning to (hyperparameter) optimize
In section 8.4 we covered connections between HPO and meta-learning from a problem-setting point of view, highlighting the connections between meta-parameters of a meta-learning algorithm and hyperparameters of a standard algorithm. In addition, one may also think of meta-learning an hyperparamter optimization algorithm, in that the task distribution is made of HPO problems, each described by a distribution over datasets and response functions of one or several standard learning algorithms. Samples from such a task distribution may involve various datasets and learning algorithms and traces of evaluations of hyperparameter configurations and their response value (e.g. validation error). The goal of the meta-learner in this context would be to learn a general strategy for efficiently searching the hyperparameter space based on data collected while optimizing hyperparameters across many different problems.
Early attempts use meta-data to warm start model-based methods, or learn loss curves; see [hutter2019automated, Chapter 2] for a survey. However, more recent works take this perspective one step further, motivated by successes of transformer-based language modelling and learning to optimize, a sub-field of meta-learning, [chen2022learning]. For example, [chen2022towards] proposed using a large language model to directly generate promising hyperparameter configurations based on a natural language description of the dataset and learning algorithm. The language model is fine-tuned on a corpus of hyperparameter optimization traces, allowing it to capture complex relationships between hyperparameters and adapt its search strategy to the problem at hand.
@@ 10.3 HPO and foundation models
Zero-shot [wei2021finetuned] and in-context learning [brown2020language] have become standard paradigms for utilizing large language models (LLM), both of which require choosing context (or prompt) templates and in-context examples. LLMs’ performances at downstream tasks may be sensitive to such choices. These choices are often carried out manually by inspecting few rounds of interactions (prompt engineering). However, both problems can be formulated as hyperparamter optimization tasks, whereby the hyperparameters may be discrete choices over in-context examples or prompt templates, or even entire sequences over tokens. The resulting search space may be large and invariably discrete, and the response functions often involve several potentially expensive and potentially noisy calls to the LLM, which render the resulting HPO problems particularly challenging. Initial attempts include AutoPrompt [shin2020autoprompt], a gradient-based approach (cf. Chapter 7) that optimizes for discrete “trigger” tokens, Self-Adaptive In-Context Learning [wu2022self], which proposes a selection-then-rank framework for selecting instance-dependent examples; TRIPLE [shi2024efficient] which tackles template optimization under a best arm identification angle (algorithms proposed include successive halving; cf. Section 5.2.1); HbBoPs [schneider2024hyperband], which leverages a structure-aware Gaussian Process surrogate combined with Hyperband to efficiently select prompts in a black-box LLM setting; Black-box Discrete Prompt Learning [diao2022black] which uses a policy gradient estimator to optimize the prompt templates, among others. Further research is needed to study more efficient and scalable methods for optimizing prompts and in-context examples for large language models. Advances in reinforcement learning and gradient-based approaches (with discrete gradient estimators – see Section 7.6) may prove useful in these directions,
@@
In addition, the emergence of agents and agent networks where various LLMs interact and act by accessing to external tools [schick2024toolformer, yang2024swe] provides another compelling application ground for HPO, in particular considering that many applications leverage closed-source LLMs queried through APIs. These agent-based systems introduce new hyperparameters to optimize, such as the selection and configuration of external tools available to the agents, including documentation, output formatting and error reporting, as well as the the communication protocols and the interactions graph between agents – see e.g. GPTSwarm [zhuge2024language]. Optimizing these hyperparameters presents unique challenges due to the complex interactions between agents, tools, and the underlying LLMs. Traditional HPO methods may need to be adapted or extended to handle the dynamic and potentially non-stationary nature of these systems. Furthermore, the use of closed-source LLMs via APIs introduces additional constraints, such as rate limits and usage costs, which must be considered in the optimization process.
@@ Acknowledgements.
The work of PF was partially supported by Italian national project CAI4DSA, CUP B13C23005640006.

### v3

@@ 8.5 Transfer of Hyperparameters Across Model Scale
In this chapter, we provide an overview of different open-source libraries and commercial systems that implement methods discussed in the previous chapters. HPO software landscape changes frequently as new problems, methods and application domains become relevant: the provided list here is by no means exhaustive and only represents a snapshot.
@@ 9.1 Open Source Libraries and Frameworks
While previous libraries, such as Spearmint, allowed to parallelize trials across processes on the same machine, more recent libraries focus more on a large-scale distributed setting to distribute the evaluation of trials across multiple instances or compute nodes. For example, RayTune [liaw2018tune], distributes the evaluation of hyperparameter configurations using the compute framework Ray. Another example is Orion, which supports both high-performance computing infrastructures and containerized environments. salinas-automl22 recently presented Syne Tune, which supports different backends that allow to distribute the HPO process either locally or on the cloud. Furthermore, it also provides a backend to simulate the evaluation of expensive objective functions using surrogate models (see Section 9.3). This allows to simulate large-scale asynchronous HPO on a single compute core.
@@
Most HPO libraries or frameworks expect the objective function either as function handle, for example in Python, or as a single training script. The search space is usually defined outside the actual training script and hyperparameters are passed as input arguments. Optuna [akiba-sigkdd19] follows a different approach by a so-called define-by-run API that allows users to dynamically define the search space instead of providing it in advance.
@@ 9.2 Commercial Systems
Since hyperparameter optimization plays such an integral role in the machine learning workflow, most machine learning platforms, for example Weights and Biases, provide some functionality to select hyperparameters. There are also several commercial systems specific for automated hyperparameter optimization. For example, Google Vizier [google_vizier] is an internal service that contributed to several Google products and research efforts. oss_vizier also introduces an open-source version of Google Vizier as a stand-alone Python package. AWS SageMaker provides a fully automated service called Automated Model Tuner [perrone2021sagemaker] for gradient-free optimization problems, which allows for large-scale HPO leveraging built-in SageMaker functionalities. Another example is Sigopt, which provides optimization-as-a-service for hyperparameter optimization utilizing Bayesian optimization techniques.
@@ 9.3 Benchmarking
Instead of using a machine learning model to approximate the true objective, klein-arxiv19a first discretized the search space of a fully-connected neural network leading to a finite number of hyperparameter configurations and then performed an exhaustive search to collect the performance and runtime of each configuration. This allows to replace the actual training and validation of a hyperparameter configuration by simple table look-ups during the HPO process. To mimic the observation noise that comes with the HPO of neural networks, each HPO configuration was evaluated multiple times. During the HPO process, random entries for each hyperparameter configuration are returned to simulate the observation noise. The idea of tabularized benchmarks has since been extended to neural architecture search [ying-icml19, dong-iclr20] (see Section 8.3) leading to several new benchmarks. While tabularized benchmarks return for each query truly observed values of the objective function instead of predictions of the model, pfisterer-automl22 argued that the discretization of the search space eliminates the local structure of the objective function, leading actually to a higher bias than using surrogate models. The idea of surrogate benchmarks has been extended to define a generative model by klein-nips19, which allows sampling an arbitrary set of new benchmarks.
To standardize the interface to different benchmarks, eggensperger_towards_2013 introduced HPOLib, which consolidates different benchmarks from the literature. In follow-up work [eggensperger-neurips21], HPOBench containerized these benchmarks to increase reproducibility of benchmarking results. HPO-B [pineda-neurips21] provides tablularized benchmarks based on OpenML datasets [vanschoren-sigkdd13a]. In a similar vein, zimmer-ieee21 introduced LCBench, which provides full learning curves of neural network models trained on OpenML datasets. salinas-automl22 also provided a unified interface to tabular and surrogate benchmarks in Syne-Tune, using a fast data format to accelerate the table look-ups. Another suite for surrogate-based benchmarking is YAHPO [pfisterer-automl22], consisting of different scenarios, including combined hyperparameter optimization and algorithm selection.
@@
By now, hyperparameter optimization may be considered a mature field with well-established algorithmic solutions and with industry-level software frameworks. It is a crucial ingredient for automated machine learning, an active research area that has even started a conference in 2022. In this monograph, we have reviewed in depth the major classes of methods (random and quasi-random search, model-based, multi-fidelity, population-based, and gradient-based) and highlighted connections to closely related areas such neural architecture search and meta-learning.
@@ 10.1 Response Functions for Unsupervised Learning
HPO requires a well-defined response function to be minimized (see Section 2.2), which in turn requires a well-defined performance metric. While suitable metrics are abundant for supervised learning, the same is not true for unsupervised learning.
@@
There have been interesting attempts to interpret this perspective in terms of precision and recall [sajjadi:2018:assessing_generative_models, simon:2019:revisiting_precision_recall], where the two supports of real and fake data are estimated and compared. Fake points off the support of real distribution are interpreted as false positives, and real points outside the support of the model’s distribution are interpreted as false negatives. The idea has sparkled a number of practically computable metrics [kynkaanniemi:2019:improved_precision_recall, naeem:2020:reliable_fidelity_diversitya, park:2023:probabilistic_precision_recall] that however have not been shown to be effective to guide HPO. A unidimensional metric that aims to assess the similarity between the model and the real data distributions is the overlapping area. Computing the overlaps between these two distributions requires reliable density estimation, which is however difficult for high dimensional data. A simple trick to overcome the associated curse of dimensionality issue is to consider instead the distribution of intra-set distances (between pairs of real data points) and inter-set distances (between pairs of fake points). This metric was introduced by yang:2020:evaluation_generative_models for assessing the quality of generative models in symbolic music, but distances in that case are based on domain specific aspects of music such as rhythm or harmony, and cannot be easily extended to other types of data. A closely related metric based on maximum mean discrepancy [obray:2021:evaluation_metrics_graph], popular in the context of graph generative models, has been shown to be complicated due to its strong sensitivity to hyperparameters in the definition of the metric itself, leading to a sort of hyper-hyperparameters.
@@
10.2 Learning to Optimize Hyperparameter
In Section 8.4, we covered connections between HPO and meta-learning from a problem-setting point of view, highlighting the connections between meta-parameters of a meta-learning algorithm and hyperparameters of a standard algorithm. In addition, one may also think of meta-learning a hyperparameter optimization algorithm, in that the task distribution is made of HPO problems, each described by a distribution over datasets and response functions of one or several standard learning algorithms. Samples from such a task distribution may involve various datasets and learning algorithms and traces of evaluations of hyperparameter configurations and their response value (e.g., the validation error). The goal of the meta-learner in this context would be to learn a general strategy for efficiently searching the hyperparameter space based on data collected while optimizing hyperparameters across many different problems.
Early attempts use meta-data to warm start model-based methods, or learn loss curves; see [hutter2019automated, Chapter 2] for a survey. However, more recent works take this perspective one step further, motivated by successes of transformer-based language modeling and learning to optimize, a sub-field of meta-learning [chen2022learning]. For example, [chen2022towards] proposed using a large language model to directly generate promising hyperparameter configurations based on a natural language description of the dataset and learning algorithm. The language model is fine-tuned on a corpus of hyperparameter optimization traces, allowing it to capture complex relationships between hyperparameters and adapt its search strategy to the problem at hand.
@@ 10.3 HPO and Foundation Models
Zero-shot [wei2021finetuned] and in-context learning [brown2020language] have become common paradigms for utilizing large language models (LLM), both of which require choosing context (or prompt) templates and in-context examples. LLMs’ performances at downstream tasks may be sensitive to such choices. These choices are often carried out manually by inspecting few rounds of interactions (prompt engineering). However, both problems can be formulated as hyperparameter optimization tasks, whereby the hyperparameters may be discrete choices over in-context examples or prompt templates, or even entire sequences over tokens. The resulting search space may be large and invariably discrete, and the response functions often involve several potentially expensive and potentially noisy calls to the LLM, which render the resulting HPO problems particularly challenging. Initial attempts include AutoPrompt [shin2020autoprompt], a gradient-based approach (see Chapter 7) that optimizes for discrete “trigger” tokens; Self-Adaptive In-Context Learning [wu2022self], which proposes a selection-then-rank framework for selecting instance-dependent examples; TRIPLE [shi2024efficient] which tackles template optimization under a best arm identification angle (algorithms proposed include successive halving; see Section 5.2.1); HbBoPs [schneider2024hyperband], which leverages a structure-aware Gaussian Process surrogate combined with Hyperband to efficiently select prompts in a black-box LLM setting; Black-box Discrete Prompt Learning [diao2022black] which uses a policy gradient estimator to optimize the prompt templates, among others. Further research is needed to study more efficient and scalable methods for optimizing prompts and in-context examples for large language models. Advances in reinforcement learning and gradient-based approaches with discrete gradient estimators (see Section 7.5.1) may prove useful in these directions.
@@
In addition, the emergence of agents and agent networks where various LLMs interact and act by accessing to external tools [schick2024toolformer, yang2024swe] provides another compelling application ground for HPO, in particular considering that many applications leverage closed-source LLMs queried through APIs. These agent-based systems introduce new hyperparameters to optimize, such as the selection and configuration of external tools available to the agents, including documentation, output formatting and error reporting, as well as the communication protocols and the interactions graph between agents — see, e.g., GPTSwarm [zhuge2024language]. Optimizing these hyperparameters presents unique challenges due to the complex interactions between agents, tools, and the underlying LLMs. Traditional HPO methods may need to be adapted or extended to handle the dynamic and potentially non-stationary nature of these systems. Furthermore, the use of closed-source LLMs via APIs introduces additional constraints, such as rate limits and usage costs, which must be considered in the optimization process.
@@ Acknowledgements.
The work of PF and MP was partially supported by the European Union NextGenerationEU and the Italian National Recovery and Resilience Plan through the Ministry of University and Research (MUR), under Projects CAI4DSA (CUP B13C23005640006) and PE0000013 (CUP J53C22003010006), respectively.